# No License

### v3

@@ 3.1 Soft-isomorphic relational graph convolution network
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})$ (5)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1 and bounded, equinumerous multisets $\bm{H}^{(1)},\bm{H}^{(2)}$. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{H}^{(1)},\bm{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
$\varepsilon_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1},$ (6)
@@
$G(\bm{H})=\sum_{\bm{h}\in\bm{H}}g(\bm{h}).$ (7)
Similarly, Corollary 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $D_{u}$ on multisets of $\mathcal{H}$ defined in Eqn. 5 with a corresponding pseudometric distance $d_{u}$ on $\mathcal{H}$, there exists a corresponding feature map $g_{u}$ and soft-injective hash function $G_{u}$ defined in Eqn. 7 that produces similar outputs for distinct multisets of neighborhood features $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ if and only if $D_{u}$ deems $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ to be sufficiently similar on some representation. Likewise, the lower and upper bounds guarantee the ability of $G_{u}$ to separate equinumerous multisets of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while maintaining the relationship with respect to $D_{u}$. In this setting, the feature map $g_{u}$ may be interpreted as the message function (Gilmer et al., 2017) of the aggregation strategy that transforms the individual neighborhood features. Meanwhile, the psuedometric $D_{u}$ corresponds to the kernel distance (Joshi et al., 2011) which intuitively represents the difference between the cross-distance and self-distance between two multisets.
@@
To illustrate the role of pseudometrics, consider anomaly detection on node features representing zero-mean scores, with node $u$ having two neighbors $v_{1}$ and $v_{2}$. If $d_{u}$ simply corresponds to the Euclidean distance, then the corresponding hash function $G_{u}$ becomes linear as presented in Fig. 2(a). The contour plot highlights collisions—instances where distinct inputs produce identical outputs (i.e., the equivalence class $\left[\bm{H}\right]_{D}$)—between dissimilar multisets of neighborhood features, resulting in aggregated neighborhood features that are less useful for the task.
@@
Nevertheless, other choices of pseudometrics, possibly incorporating prior knowledge, would correspond to more complex message functions $g_{u}$. This leads to non-trivial hash functions $G_{u}$ and contour plots where only the regions determined by $D_{u}$ to be similar may produce similar aggregated neighborhood features, making collisions more informative and controlled. This also highlights the significance of dynamic (Brody et al., 2021) or non-linear message functions $g_{u}$ in MPNNs.
As further illustration, if $d_{u}$ instead corresponds to the Euclidean distance of the squared score, then the corresponding hash function $G_{u}$ has the contour plot in Fig. 2(b). The resulting hash collisions and equivalence classes then become more useful and meaningful for detecting anomalous scores.
@@
$D^{2}\left(\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)};\bm{h_{u}}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(1)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(2)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}}),$ (8)
@@
$G\left(\bm{H_{u}};\bm{h_{u}}\right)=\sum_{\bm{h}\in\bm{H_{u}}}g\left(\bm{h};\bm{h_{u}}\right),$ (9)
for every node $u\in\mathcal{V}$. This approach makes $D$, $d$, $g$, and $G$ anisotropic (Dwivedi et al., 2023) (i.e., a function of both the features of the query (center) node $\bm{h_{u}}$ and key (neighboring) nodes $\bm{h}\in\bm{H_{u}}$). Moreover, contextualized on the features of the query node, $D$ may still be interpreted as a pseudometric controlling hash collisions with a corresponding soft-injective hash function $G$.
Furthermore, the integration of $\bm{h_{u}}$ also allows for the interpretation of $g$ as a relational message function, guiding how features of the key nodes are to be embedded and transformed based on the features of the query node. Figs. 2(c) and 2(d) provide intuition for this idea where the introduction of a bias term, assuming a function of the features of the query node, shifts the contour plot to produce distinct aggregated neighborhood features $\bm{a_{u}}\neq\bm{a_{u^{\prime}}}$ for nodes $u$ and $u^{\prime}$ with identical neighborhood features $\bm{H_{u}}=\bm{H_{u^{\prime}}}$ but distinct features $\bm{h_{u}}\neq\bm{h_{u^{\prime}}}$. Nevertheless, one may also inject stochasticity into the node features to distinguish between nodes $u$ and $u^{\prime}$ with identical features $\bm{h_{u}}=\bm{h_{u^{\prime}}}$ and neighborhood features $\bm{H_{u}}=\bm{H_{u^{\prime}}}$ with high probability (Sato et al., 2021) and to imitate having distinct $D_{u}$, $d_{u}$, $g_{u}$, and $G_{u}$ for each node $u\in\mathcal{V}$.
@@
For a graph representation learning problem, one may directly model the anisotropic and dynamic relational message function $g$ as a two-layer multi-layer perceptron (MLP), with implicitly learned pseudometrics, following the universal approximation theorem (Hornik et al., 1989) to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
$\bm{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (10)
where $\sigma$ is a non-linear activation function, $\bm{W_{Q}},\bm{W_{K}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$, and $\bm{W_{R}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{hidden}}}$. Leveraging linearity, the model has a computational complexity of
@@
with computational efficiency achieved by the application of only an activation function along edges, making it comparable to GAT. Nevertheless, $\sigma$ may also be replaced with a deep MLP in practice if modeling $g$ as a shallow two-layer MLP becomes infeasible.
In essence, the proposed SIR-GCN is a simple, interpretable, and computationally efficient instance of the MPNN framework. Moreover, in contrast to other MPNN instances in literature, the proposed model emphasizes the anisotropic and dynamic transformation of neighborhood features to obtain contextualized messages.
@@ 3.2 Soft-isomorphic graph readout function
Corollary 1 also shows that, for each graph $\mathcal{G}$, given a pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn. 5, there exists a corresponding feature map $r_{\mathcal{G}}$ and soft-injective graph readout function $R_{\mathcal{G}}$ defined in Eqn. 7. While this result holds for each graph $\mathcal{G}$ independently, one may simply consider a single $D$, $d$, $r$, and $R$ for every graph $\left\{\mathcal{G}_{d}\right\}_{d\in\mathcal{D}}$ under task $\mathcal{D}$. Nevertheless, the graph context and structure may also be integrated into $D$, $d$, $r$, and $R$, through a virtual super node (Gilmer et al., 2017) for instance, to imitate having distinct $D_{\mathcal{G}}$, $d_{\mathcal{G}}$, $r_{\mathcal{G}}$, and $R_{\mathcal{G}}$ for each graph $\mathcal{G}$ and to further enhance its representational capability.
Similarly, for a graph representation learning problem, $r$ may also be directly modeled as an MLP, with implicitly learned pseudometrics, to obtain the soft-isomorphic graph readout function
$\bm{h_{\mathcal{G}}}=\sum_{v\in\mathcal{V}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\bm{h_{v}}\right),$ (12)
where $\textsc{MLP}_{R}$ corresponds to $r$ and $\bm{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
@@ 4 Mathematical discussion
The mathematical relationship of SIR-GCN with GCN, GraphSAGE, GAT, GIN, and PNA are presented in this section to highlight the novelty and contribution. While activation functions and MLPs applied after each GNN layer play a significant role in the overall performance, the discussions only focus on the message-passing operation that defines GNNs. The relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the former.
@@ 4.1 GCN and GraphSAGE
It may be shown that Corollary 1 holds up to a constant scale. Hence, the mean aggregation and symmetric mean aggregation, by extension, may be used in place of the sum aggregation. If one sets $\sigma$ as identity or $\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, $\bm{W_{R}}\bm{W_{K}}=\bm{W}$, and $\tilde{\mathcal{N}}(u)=\mathcal{N}(u)\cup\{u\}$, one obtains
$\bm{h^{*}_{u}}=\sum_{v\in\mathcal{N}(u)}\dfrac{1}{\sqrt{\left|\mathcal{N}(u)\right|}\sqrt{\left|\mathcal{N}(v)\right|}}\bm{W}\bm{h_{v}}$ (13)
@@
$\bm{h^{*}_{u}}=\dfrac{1}{\left|\tilde{\mathcal{N}}(u)\right|}\sum_{v\in\tilde{\mathcal{N}}(u)}\bm{W}\bm{h_{v}}$ (14)
which recovers GCN and GraphSAGE with mean aggregation, respectively. Moreover, the sum aggregation may also be replaced with the max aggregation, albeit without theoretical justification, to recover GraphSAGE with max pooling. Thus, GCN and GraphSAGE may be viewed as instances of SIR-GCN. [^2] GraphSAGE with LSTM aggregation is not included in this discussion. The difference lies in the isotropic (Dwivedi et al., 2023) nature (i.e., a function of only the features of the key nodes) of GCN and GraphSAGE and the use of non-linearities only in the combination strategy.

### v4

@@ 3.1 Soft-isomorphic relational graph convolution network
$D^{2}\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)\coloneq\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(1)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(2)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})$ (6)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1 and bounded, equinumerous multisets $\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}$. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
$\varepsilon_{2}<\left\lVert G\left(\boldsymbol{H}^{(1)}\right)-G\left(\boldsymbol{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)<\varepsilon_{1},$ (7)
@@
$G(\boldsymbol{H})=\sum_{\boldsymbol{h}\in\boldsymbol{H}}g(\boldsymbol{h}).$ (8)
Similarly, Corollary 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $D_{u}$ on multisets of $\mathcal{H}$ defined in Eqn. 6 with a corresponding pseudometric distance $d_{u}$ on $\mathcal{H}$, there exists a corresponding feature map $g_{u}$ and hash function $G_{u}$ defined in Eqn. 8 that produces similar outputs for distinct multisets of neighborhood features $\boldsymbol{H_{u}}^{(1)},\boldsymbol{H_{u}}^{(2)}$ if and only if $D_{u}$ deems $\boldsymbol{H_{u}}^{(1)}$ and $\boldsymbol{H_{u}}^{(2)}$ to be sufficiently similar on some representation. Likewise, the lower and upper bounds guarantee the ability of $G_{u}$ to separate equinumerous multisets of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while maintaining the relationship with respect to $D_{u}$. In this setting, the feature map $g_{u}$ may be interpreted as the soft-injective message function (Gilmer et al., 2017) that transforms the individual neighborhood features with a corresponding soft-injective hash function $G_{u}$. Meanwhile, the pseudometric $D_{u}$ corresponds to the kernel distance (Joshi et al., 2011) which intuitively represents the difference between the cross-distance and self-distance between two multisets. The two necessary properties of the soft-injective message function—dynamic and anisotropic—are then motivated below.
@@
To illustrate the role of the pseudometric, consider node $u$ with two neighbors $v_{1}$ and $v_{2}$ and the task of anomaly detection on the scalar node features $\boldsymbol{h_{v_{1}}}$ and $\boldsymbol{h_{v_{2}}}$ representing zero-mean scores. If $d_{u}$ simply corresponds to the Euclidean distance, then the corresponding hash function $G_{u}$ becomes linear as seen in Fig. 2(a). Crucially, the contour plot highlights collisions—instances where distinct inputs produce identical outputs (i.e., the equivalence class $\left[\boldsymbol{H}\right]_{D}$)—between dissimilar multisets of neighborhood features, resulting in aggregated neighborhood features that are less useful for the task.
@@
Nevertheless, other choices of pseudometrics, possibly incorporating prior knowledge, would correspond to more complex message functions $g_{u}$. This leads to non-trivial hash functions $G_{u}$ and contour plots where only the regions determined by $D_{u}$ to be similar with respect to a given task may produce similar aggregated neighborhood features, making collisions more informative and controlled. In practice, this also highlights the significance of dynamic (Brody et al., 2022) (i.e., a universal function approximator (Hornik et al., 1989)) message functions $g_{u}$ in the MPNN framework, which may be modeled as multi-layer perceptrons (MLPs) as illustrated in Figs. 2(b) and 2(c).
As further illustration, if $d_{u}$ instead corresponds to the Euclidean distance of the squared score, then the corresponding hash function $G_{u}$ has the contour plot in Fig. 2(d). The resulting hash collisions and equivalence classes then become more useful and meaningful for detecting anomalous scores.
@@
$D^{2}\left(\boldsymbol{H_{u}}^{(1)},\boldsymbol{H_{u}}^{(2)};\boldsymbol{h_{u}}\right)\coloneq\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H_{u}}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}};\boldsymbol{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H_{u}}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H_{u}}^{(1)}\end{subarray}}\!\!d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}};\boldsymbol{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H_{u}}^{(2)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}};\boldsymbol{h_{u}}),$ (9)
@@
$G\left(\boldsymbol{H_{u}};\boldsymbol{h_{u}}\right)\coloneq\sum_{\boldsymbol{h}\in\boldsymbol{H_{u}}}g\left(\boldsymbol{h};\boldsymbol{h_{u}}\right),$ (10)
for every node $u\in\mathcal{V}$. This approach makes $D$, $d$, $g$, and $G$ anisotropic (Dwivedi et al., 2023) (i.e., a function of both the features of the query (center) node $\boldsymbol{h_{u}}$ and key (neighboring) nodes $\boldsymbol{h}\in\boldsymbol{H_{u}}$). In addition, contextualized on the features of the query node, $D$ may still be interpreted as a pseudometric controlling hash collisions with a corresponding soft-injective hash function $G$.
Furthermore, the integration of $\boldsymbol{h_{u}}$ also allows for the interpretation of $g$ as a soft-injective relational message function, guiding how features of the key nodes are to be embedded and transformed based on the features of the query node. Figs. 2(b) and 2(c) provide intuition for this idea where the introduction of a bias term, assuming a function of the features of the query node, shifts the contour plot to produce distinct aggregated neighborhood features $\boldsymbol{a_{u}}\neq\boldsymbol{a_{u^{\prime}}}$ for nodes $u\neq u^{\prime}\in\mathcal{V}$ with identical neighborhood features $\boldsymbol{H_{u}}=\boldsymbol{H_{u^{\prime}}}$ but distinct features $\boldsymbol{h_{u}}\neq\boldsymbol{h_{u^{\prime}}}$. Moreover, one may also inject stochasticity into the node features to distinguish between nodes $u\neq u^{\prime}\in\mathcal{V}$ with identical features $\boldsymbol{h_{u}}=\boldsymbol{h_{u^{\prime}}}$ and neighborhood features $\boldsymbol{H_{u}}=\boldsymbol{H_{u^{\prime}}}$ with high probability (Sato et al., 2021) and to imitate having distinct $D_{u}$, $d_{u}$, $g_{u}$, and $G_{u}$ for each node $u\in\mathcal{V}$.
@@
For a graph representation learning problem, one may directly model the anisotropic and dynamic soft-injective relational message function $g$ as a two-layer MLP, with implicitly learned pseudometrics, to obtain the soft-isomorphic relational graph convolution network (SIR-GCN)
$\boldsymbol{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\boldsymbol{W_{R}}~{}\sigma\left(\boldsymbol{W_{Q}}\boldsymbol{h_{u}}+\boldsymbol{W_{K}}\boldsymbol{h_{v}}\right),$ (11)
where $\sigma$ is a non-linear activation function, $\boldsymbol{W_{Q}},\boldsymbol{W_{K}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{in}}}$, and $\boldsymbol{W_{R}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{hidden}}}$. Leveraging linearity, the model has a computational complexity of
@@
with only the application of an activation function along edges, making it comparable to classical GNNs in literature. Nevertheless, $\sigma$ may also be modeled as a deep MLP in practice if modeling $g$ as a shallow two-layer MLP becomes infeasible.
In essence, SIR-GCN is a simple instance of the MPNN framework explicitly designed to handle uncountable node features while maintaining rigorous theoretical foundations. It emphasizes the anisotropic and dynamic transformation of neighborhood features, obtaining contextualized messages that enable it to learn complex relationships between neighboring nodes. Moreover, the proposed model is also computationally efficient, requiring only a single aggregator and applying only an activation function along edges to facilitate effective message-passing of uncountable node features.
@@ 3.2 Soft-isomorphic graph readout function
Corollary 1 also shows that, for each graph $\mathcal{G}$, given a pseudometric distance $D_{\mathcal{G}}$ on multisets of $\mathcal{H}$ defined in Eqn. 6 with a corresponding pseudometric distance $d_{\mathcal{G}}$ on $\mathcal{H}$, there exists a corresponding feature map $r_{\mathcal{G}}$ and graph readout function $R_{\mathcal{G}}$ defined in Eqn. 8. While this result holds for each graph $\mathcal{G}$ independently, one may simply consider a single $D$, $d$, $r$, and $R$ for every graph $\left\{\mathcal{G}_{d}\right\}_{d\in\mathcal{D}}$ under task $\mathcal{D}$. Nevertheless, the graph context and structure may also be integrated into $D$, $d$, $r$, and $R$, through a virtual super node (Gilmer et al., 2017) for instance, to imitate having distinct $D_{\mathcal{G}}$, $d_{\mathcal{G}}$, $r_{\mathcal{G}}$, and $R_{\mathcal{G}}$ for each graph $\mathcal{G}$ and to further enhance its representational capability.
Similarly, for a graph representation learning problem, the dynamic soft-injective feature map $r$ may also be directly modeled as an MLP, with implicitly learned pseudometrics, to obtain the soft-isomorphic graph readout function
$\boldsymbol{h_{\mathcal{G}}}=\sum_{v\in\mathcal{V}_{\mathcal{G}}}\textsc{MLP}_{R}\left(\boldsymbol{h_{v}}\right),$ (13)
where $\textsc{MLP}_{R}$ corresponds to $r$ and $\boldsymbol{h_{\mathcal{G}}}$ is the graph-level feature of graph $\mathcal{G}$.
@@ 4 Mathematical discussion
The mathematical relationship between SIR-GCN and key GNNs in literature—GCN, GraphSAGE, GAT, GIN, and PNA—are presented in this section to underscore the contribution and distinctive advantages of the proposed model. It is worth noting that while activation functions and MLPs applied after each GNN layer play a significant role in the overall performance, the discussions only focus on the core message-passing operation that defines GNNs. In addition, the relationship between SIR-GCN and the 1-WL test is also presented to contextualize the representational capability of the former.
@@ 4.1 GCN and GraphSAGE
It may be shown that Corollary 1 holds up to a constant scale. Hence, the mean aggregation and symmetric mean aggregation, by extension, may be used in place of the sum aggregation in Eqn. 11. If one sets $\sigma$ as identity or $\textsc{PReLU}(\alpha=1)$, $\boldsymbol{W_{Q}}=\boldsymbol{0}$, $\boldsymbol{W_{R}}\boldsymbol{W_{K}}=\boldsymbol{W}$, and $\tilde{\mathcal{N}}(u)\coloneq\mathcal{N}(u)\cup\{u\}$, one obtains
$\boldsymbol{h^{*}_{u}}=\sum_{v\in\mathcal{N}(u)}\dfrac{1}{\sqrt{\left|\mathcal{N}(u)\right|}\sqrt{\left|\mathcal{N}(v)\right|}}\boldsymbol{W}\boldsymbol{h_{v}}$ (14)
@@
$\boldsymbol{h^{*}_{u}}=\dfrac{1}{\left|\tilde{\mathcal{N}}(u)\right|}\sum_{v\in\tilde{\mathcal{N}}(u)}\boldsymbol{W}\boldsymbol{h_{v}}$ (15)
which are equivalent to GCN and GraphSAGE with mean aggregation, respectively. Moreover, the sum aggregation may also be replaced with the max aggregation, albeit without theoretical justification, to recover GraphSAGE with max pooling. Thus, GCN and GraphSAGE [^2] GraphSAGE with LSTM aggregation is not included in this discussion. may be viewed as instances of SIR-GCN. The key difference lies in the isotropic (Dwivedi et al., 2023) nature (i.e., a function of only the features of the key nodes) of GCN and GraphSAGE and their non-linearities only after aggregating the neighborhood features.